OpenStudy (anonymous):

a relay race is 1/10 kilometers long. four athletes will run an equal distance to complete the relay. how far does each athlete run?

OpenStudy (anonymous):

so lets say that x is the number of kilometer that each athlete ran in this race. since we have 4 runners that means we have 4 x's. when we add them up we have x+x+x+x now whatever this is added to be must equal 1/10 kilometers but for now lets leave the kilometers part out. x+x+x+x =1/10 now the left side can be simplified to 4x=1/10 now we want to solve for x. All have to do now is divide both sides by 4 and you are done.